Convert to a percent.
42.089

Found 2 solutions by ewatrrr, cinna:
Answer by ewatrrr(24785)

 
Hi
Convert 42.089 to a percent by multiplying by 100
42.089*100 = 4208.9 %
CHECKING our Answer***
4208.9% = 4208.9/100 = 42.089

Answer by cinna(4)





Convert 42.089 to a percent:

1) Multiply by 100
42.089*100 = 4208.9%


Answer check:
4208.9% = 4208.9/100 = 42.089